Find the minimum and the maximum values of 122 + (p + 1)i| on the closed disc {z € C: |Z| <q+1}.Find the minimum and the maximum values of |z2 + (p + 1)i| on the closed disc {z ∈ C : |z| ≤ q + 1}
Answer #1

Que> & also (2²+ (p+1); 1 = 12²/+ | P+1);/ using triangle inequality /2/5q+1 Izimax + (R+1) = (2+1)²+ (P+1) 12²+ (p+1); max [